OpenStudy (anonymous):

Factor completely: x2 − 3x − 28 (x − 7)(x + 4) (x − 4)(x + 7) (x + 28)(x − 1) (x − 14)(x + 2)

OpenStudy (anonymous):

x^2 -3x -28 = 0 (x -7)(x + 4) = 0 x - 7 = 0 => x = 7 or x + 4 = 0 => x = -4